# CRABS: A syntactic-semantic pincer strategy for bounding LLM interpretation of Python notebooks

### v1

@@ Abstract
Recognizing the information flows and operations comprising data science and machine learning Python notebooks is critical for evaluating, reusing, and adapting notebooks for new tasks. Investigating a notebook via re-execution often is impractical due to the challenges of resolving data and software dependencies. While Large Language Models (LLMs) pre-trained on large codebases have demonstrated effectiveness in understanding code without running it, we observe that they fail to understand some realistic notebooks due to hallucinations and long-context challenges. To address these issues, we propose a notebook understanding task yielding an information flow graph and corresponding cell execution dependency graph for a notebook, and demonstrate the effectiveness of a pincer strategy that uses limited syntactic analysis to assist full comprehension of the notebook using an LLM. Our Capture and Resolve Assisted Bounding Strategy (CRABS) employs shallow syntactic parsing and analysis of the abstract syntax tree (AST) to capture the correct interpretation of a notebook between lower and upper estimates of the inter-cell I/O sets, then uses an LLM to resolve remaining ambiguities via cell-by-cell zero-shot learning, thereby identifying the true data inputs and outputs of each cell. We evaluate and demonstrate the effectiveness of our approach using an annotated dataset of 50 representative, highly up-voted Kaggle notebooks that together represent 3454 actual cell inputs and outputs. The LLM correctly resolves 1397 of 1425 (98%) ambiguities left by analyzing the syntactic structure of these notebooks. Across 50 notebooks, CRABS achieves average F1 scores of 98% identifying cell-to-cell information flows and 99% identifying transitive cell execution dependencies. Furthermore, 37 out of the 50 (74%) individual information flow graphs and 41 out of 50 (82%) cell execution dependency graphs match the ground truth exactly.
@@ 1 Introduction
Computational notebooks, such as Jupyter notebooks (Kluyver et al., 2016), provide a readable and executable medium combining code, results, and explanations that facilitates sharing of notebook functionality among users. Generally, users find it necessary to execute a notebook to discover how it works in detail, confirm that it works correctly, and assess its potential for reuse. However, only one in four [^1] Among 1,081,702 unique Python notebooks discovered on GitHub, Pimentel et al. (2019) attempted to execute the 863,878 notebooks for which both Python version and execution order could be determined, and found that only 24% of the latter notebooks could be executed without errors using either the declared dependencies or a comprehensive, preconfigured Anaconda environment. notebooks discovered on GitHub that specify both Python version and execution order can be re-executed without errors (Pimentel et al., 2019). Most of these errors stem from issues related to module dependencies and data accessibility which are often difficult or even impossible to resolve. For this reason, facilitating the interpretability of notebooks in the absence of re-exeution would represent a major step toward improving reproducibility and lowering barriers to reuse. Such interpretability requires understanding what individual variables mean, where they are defined, and where they are subsequently used. Answers to such questions are critical for evaluating, reusing, and adapting a notebook for new tasks.
@@
Extracting the information flow graph from a computational notebook requires not just identifying how data flows between code cells, but also how user-defined functions and classes declared in one cell are used another. Fortunately, recognizing code cells that define or call user-defined functions and classes is relatively straightforward via syntactic parsing and analysis of the abstract syntax tree (AST). In contrast, identifying which cells actually assign new values to variables or update them is challenging without comprehensively analyzing the source code of imported modules or executing the cells. For example, the statement phone_data.dropna(inplace=True) updates the phone_data object in place, while the statement phone_data.tuncate(before=1, after=3) returns information from the phone_data object without modifying it. Without understanding the mechanisms of these two methods, or comparing the phone_data object before and after executing the code, it is not possible to determine whether the phone_data object is updated or not. Therefore, how data flows between code cells can be ambiguous if only the syntactic structure apparent in code cells is taken into account.
@@
We observe that experienced data scientists often can resolve these ambiguities without accessing source code for imported modules and without executing the cells by hand. This raises the question: can Large Language Models (LLMs) do the same thing? LLMs pre-trained on large codebases have shown to be effective on code-related tasks, e.g., bug detection (Wang et al., 2024), code understanding (Nam et al., 2024), code completion (Zhang et al., 2023; Cheng et al., 2024), code generation (Yin et al., 2023), and code summarization (Haldar & Hockenmaier, 2024). For these reasons, we hypothesize that LLMs can help resolve ambiguites left by purely analyzing the syntactic structure of the notebook itself, i.e., just the code in its cells.
@@
Figure 1: An overview of CRABS using the example notebook in Figure 4. CRABS first produces lower and upper estimates of inter-cell I/O sets in a syntactic phase, then resolves the differences between these estimates using an LLM in a semantic-aware phase, and finally constructs a information flow graph where thick blue edges indicating these resolved sets.
@@
At present, however, LLMs cannot reliably analyze entire notebooks on their own. As we show in Section 5.3, LLMs face hallucination issues and long-context challenges with the result that they sometimes identify variables that do not exist and for longer notebooks frequently return information flow graphs with the wrong number of code cells. To address these issues, we adopt an approach that reduces the overall workload assigned to the LLM and limits its tasks to those that entail semantically informed analysis of the code.
@@
In this paper, we propose the Capture and Resolve Assisted Bounding Strategy (CRABS), a syntactic-semantic hybrid strategy for extracting the information flow graph from a Python notebook. Figure 1 is an overview of the steps employed by CRABS to generate an information flow graph for a given Python notebook using CRABS. We first produce lower and upper estimates of inter-cell I/O sets based on its syntactic structure, and then prompt an LLM to resolve the differences between these two estimates (i.e., ambiguities) one by one. To evaluate this approach, we apply CRABS to 50 representative, highly up-voted Python notebooks from the Meta Kaggle Code dataset (Plotts & Risdal, 2023).
@@ 2 Related Work
LLMs and computational notebooks. Pre-trained language models have been leveraged to generate feedback on answers to open-ended questions about instructional coding or math notebooks (Matelsky et al., 2023); and fine-tuned for code generation (Huang et al., 2022; Yin et al., 2023; Li et al., 2023; Huang et al., 2024) and code edits (Jin et al., 2025). Cao et al. (2024) explore multimodal agents for data science notebook generation.
@@ 3 A Model of Information Flow in Python Notebooks
CRABS views a Python notebook as a sequence of $n$ code cells $C=(c_{1},c_{2},...,c_{n})$ where $c_{k}$ is the $k$th code cell (markdown cells are ignored). CRABS further assumes that cells are executed in order, i.e., from the top to the bottom of the notebook. This assumption limits the possible interpretations of information flows in a notebook; i.e., the value assigned to a variable in cell $c_{k}$ cannot be used in cell $c_{k-1}$ or earlier. We represent the distinct units of information (i.e., variable values and code references) that flow between cells as $I=\{i_{k}\}_{1}^{m}$ and define an information flow graph as a set of information flows $F\subseteq\{(c_{s},c_{t},i_{k})\mid 1\leq s<t\leq n\}$, where $i_{k}$ flows from source cell $c_{s}$ to target cell $c_{t}$. The information flow graph in turn implies a cell execution dependency graph, i.e., the set of transitive dependencies $D\subseteq\{(c_{t},c_{s})\mid 1\leq s<t\leq n\}$, where cell $c_{t}$ directly or indirectly depends on cell $c_{s}$. The cell execution dependency graph can be derived from the information flow graph by (1) recognizing a direct dependency between two cells if at least one information flow exists between them in the information flow graph; and (2) computing the transitive closure of this graph to infer indirect dependencies.
@@ 4 Elements of CRABS
CRABS is based on the observation that some inter-cell I/O sets are certain by shallow syntactic parsing and analysis of the AST, while some are uncertain but may be resolved by taking the semantic meaning of these code cells into consideration. These inter-cell I/O sets are critical for further extraction of information flows and transitive dependencies to construct their corresponding graphs. CRABS performs the notebook understanding task in two phases: a syntactic phase and a semantic-aware phase. The syntactic phase employs shallow syntactic parsing to extract an AST based purely on the notebook itself, and analyzes the extracted AST to yield lower and upper estimates of the inter-cell I/O sets. The semantic-aware phase resolves the ambiguities between these lower and upper estimates using an LLM.
We represent inter-cell I/O sets as $S\subset\{(c,i_{k},\textit{tag})\mid c\in\{c_{s},c_{t}\},\textit{tag}\in\{\texttt{out},\texttt{in}\}\}$, where the out tag denotes that information $i_{k}$ flows out of source cell $c_{s}$, and the in tag denotes that information $i_{k}$ flows into target cell $c_{t}$. Therefore, it is critical to identify four items for each cell, i.e., data inputs, data outputs, code declarations, and code references. If a data $i_{k}$ is defined in cell $c_{s}$ and used in cell $c_{t}$, then $i_{k}$ is a data output of $c_{s}$ and a data input of $c_{t}$. However, when cells are examined in isolation, we cannot determine whether $i_{k}$ is used in $c_{t}$ or not when only $c_{s}$ is given. Because we do not provide any information about subsequent cells when asking the LLM to resolve ambiguous information flows associated with a particular cell, we refer to data output candidates rather than data outputs in these prompts. Code includes user-defined functions and classes, and are handled analogously to data. In the rest of this section, we describe syntactic phase and semantic-aware phase based on these four items. Relevant examples are attached in the Appendix.
@@ 4.1 Syntactic Phase
The syntactic phase of CRABS performs a limited static analysis of the code in a notebook to yield two estimates of the inter-cell I/O sets, representing lower and upper estimates of the actual flows of information into or out of each cell. It produces two estimates because CRABS aims to interpret Python notebooks without accessing the source code for any dependencies, e.g., code defined outside the notebook and accessed via function calls. Deeper or more sophisticated syntactic parsing and analysis of the AST that descends into each function called from a cell likely would achieve better lower and upper estimates of the inter-cell I/O sets, leaving fewer or even no ambiguities to be resolved by an LLM. However, because we restrict our syntactic analysis to code explicitly included in a notebook’s cells, CRABS does not need to resolve software dependencies. Moreover, even comprehensive static analysis will not yield a unique prediction of the dataflow between cells when conditional control flow structures are employed, e.g., if variables are assigned, used, or updated conditionally (see Figure 10 in the Appendix for an example).
@@
CRABS makes three basic assumptions about the code in Python notebooks relevant to the syntactic phase. First, execution flow follows a strict top-to-bottom order, starting from an initial state with all variables, imports, and other runtime information cleared from memory. Second, all code is reachable. And third, function and class names are not reused as variable names or vice versa (see Appendix A.2.1). In addition, our current implementation has three further limitations. Global variables that are accessed or modified within a function, without being explicitly passed as arguments, will not be recognized as data inputs or data output candidates by our system on this basis (see Appendix A.2.2). Lines of code employing languages other than Python, e.g., shell commands invoked using the ! syntax, or IPython magic commands invoked using the % syntax are not taken into account (see Appendix A.2.3). A final limitation related to shared data references (aliased variables) is more complex. This issue arises when multiple variables reference the same mutable object in memory, and one of them is modified in place. For example, the first code cell in Figure 5 defines two Pandas DataFrames, train and test. The datasets object is a list that references two mutable objects, train and test. In the second code cell, the train is modified in place which results in a change in datasets. We define this phenomenon as hidden modifications, wherein a variable name (i.e., datasets) is not explicitly written in a code cell but should still be considered a data output candidate. The current implementation of CRABS detects only those shared references explicitly indicated by assignment (e.g., x = y) or by inclusion in collections (i.e., mutable objects stored in a list, a tuple, or a dictionary). This covers all instances of shared references in our dataset; we will extend CRABS to detect shared references established by other mechanisims in the future.
@@
The lower estimate computed by the syntactic phase represents the set of cell inputs and output candidates that are certain (i.e., unambiguous) given just the code apparent in the cell. This lower estimate is computed using the following rules: (1) Variables explicitly passed to a function are regarded as data inputs only. (2) The iterator of a for loop, a while loop, or a comprehension is regarded as a data input only, unless it is directly reassigned in the loop body. (3) Code on all branches of a conditional statement are considered to contribute to the inputs and the output candidates of the cell. (4) When a conditional statement occurs inside a loop, the intersection of all possible execution orders are considered (see Appendix A.2.4 for an illustrative example). (5) Shared references (i.e., aliasing) are not considered. Figure 9 and Figure 9 show the information flow graph and the corresponding cell execution dependency graph generated from the lower estimate.
The upper estimate represents the set of cell inputs and output candidates that are either certain and ambiguous, which is a superset of both the ground truth and the lower estimate. This upper estimate is computed using the following rules: (1) Variables explicitly passed to a function are regarded as both data inputs and data output candidates. (2) The iterator of a for loop, a while loop, or a comprehension is regarded both as a data input and a data output candidate. (3) As with the lower estimate, code on all branches of a conditional statement are considered to contribute to the inputs and the output candidates of the cell. (4) When a conditional statement occurs inside a loop, the union of all possible execution orders are considered. (5) Shared references are considered. Figure 9 and Figure 9 show the information flow graph and the corresponding cell execution dependency graph generated from the upper estimate.
Figure 6: An information flow graph generated from the lower estimate of the example notebook in Figure 4.
Figure 7: A cell execution dependency graph extracted from Figure 9.
Figure 8: An information flow graph generated from the upper estimate of the example notebook in Figure 4.
Figure 9: A cell execution dependency graph extracted from Figure 9.
Figure 9 and  9 demonstrate a “subgraph/supergraph” relation between the cell execution dependency graphs generated from lower and upper estimates, and the corresponding ground truth, i.e., the graph generated from the lower estimate is a subgraph of the ground truth, and the graph generated from the upper estimate is a supergraph of the ground truth.

### v2

@@ Abstract
Recognizing the information flows and operations comprising data science and machine learning Python notebooks is critical for evaluating, reusing, and adapting notebooks for new tasks. Investigating a notebook via re-execution often is impractical due to the challenges of resolving data and software dependencies. While Large Language Models (LLMs) pre-trained on large codebases have demonstrated effectiveness in understanding code without running it, we observe that they fail to understand some realistic notebooks due to hallucinations and long-context challenges. To address these issues, we propose a notebook understanding task yielding an information flow graph and corresponding cell execution dependency graph for a notebook, and demonstrate the effectiveness of a pincer strategy that uses limited syntactic analysis to assist full comprehension of the notebook using an LLM. Our Capture and Resolve Assisted Bounding Strategy (CRABS) employs shallow syntactic parsing and analysis of the abstract syntax tree (AST) to capture the correct interpretation of a notebook between lower and upper estimates of the inter-cell I/O set—the flows of information into or out of cells via variables—then uses an LLM to resolve remaining ambiguities via cell-by-cell zero-shot learning, thereby identifying the true data inputs and outputs of each cell. We evaluate and demonstrate the effectiveness of our approach using an annotated dataset of 50 representative, highly up-voted Kaggle notebooks that together represent 3454 actual cell inputs and outputs. The LLM correctly resolves 1397 of 1425 (98%) ambiguities left by analyzing the syntactic structure of these notebooks. Across 50 notebooks, CRABS achieves average F1 scores of 98% identifying cell-to-cell information flows and 99% identifying transitive cell execution dependencies. Moreover, 37 out of the 50 (74%) individual information flow graphs and 41 out of 50 (82%) cell execution dependency graphs match the ground truth exactly.
@@ 1 Introduction
Computational notebooks, such as Jupyter notebooks (Kluyver et al., 2016), provide a readable and executable medium combining code, results, and explanations that facilitates sharing of notebook functionality among users. Generally, users find it necessary to execute a notebook to discover how it works in detail, confirm that it works correctly, and assess its potential for reuse. However, only one in four [^1] Among 1,081,702 unique Python notebooks discovered on GitHub, Pimentel et al. (2019) attempted to execute the 863,878 notebooks for which both Python version and execution order could be determined, and found that only 24% of the latter notebooks could be executed without errors using either the declared dependencies or a comprehensive, preconfigured Anaconda environment. notebooks discovered on GitHub that specify both Python version and execution order can be re-executed without errors (Pimentel et al., 2019). Most of these errors stem from module dependency and data accessibility issues that are often difficult or even impossible to resolve. For this reason, facilitating interpretability of notebooks in the absence of re-exeution would represent a major step toward improving reproducibility and lowering barriers to reuse. Such interpretability requires understanding what individual variables mean, where they are defined, and where they are subsequently used. Answers to such questions are critical for evaluating, reusing, and adapting a notebook for new tasks.
@@
Extracting the information flow graph from a computational notebook requires not just identifying how data flows between code cells, but also how user-defined functions and classes declared in one cell are used another. Fortunately, recognizing code cells that define or call user-defined functions and classes is relatively straightforward via syntactic parsing and analysis of the abstract syntax tree (AST). In contrast, identifying which cells actually assign new values to variables or update them is challenging without comprehensively analyzing the source code of imported modules or executing the cells. For example, the statement phone_data.dropna(inplace=True) updates the phone_data object in place, while the statement phone_data.truncate(before=1, after=3) returns information from the phone_data object without modifying it. Without understanding the mechanisms of these two methods, or comparing the phone_data object before and after executing the code, it is not possible to determine whether the phone_data object is updated or not. Therefore, how data flows between code cells can be ambiguous if only the syntactic structure apparent in code cells is taken into account.
@@
We observe that experienced data scientists often can resolve these ambiguities without accessing source code for imported modules and without executing the cells by hand. This raises the question: can Large Language Models (LLMs) do the same thing? LLMs pre-trained on large codebases have been shown to be effective on code-related tasks, e.g., bug detection (Wang et al., 2024), code understanding (Nam et al., 2024), code completion (Zhang et al., 2023; Cheng et al., 2024), code generation (Yin et al., 2023), and code summarization (Haldar & Hockenmaier, 2024). For these reasons, we hypothesize that LLMs can help resolve ambiguites left by purely analyzing the syntactic structure of the notebook itself, i.e., just the code in its cells.
@@
Figure 1: An overview of CRABS using the example notebook in Figure 4. CRABS first produces lower and upper estimates of the inter-cell I/O set in a syntactic phase, then resolves the differences between these estimates using an LLM in a semantic-aware phase, and finally constructs an information flow graph with thick blue edges indicating this resolved set.
@@
At present, however, LLMs cannot reliably analyze entire notebooks on their own. As we show in Section 5.3, LLMs face hallucination issues and long-context challenges with the result that they sometimes identify variables that do not exist, and for longer notebooks frequently return information flow graphs with the wrong number of code cells. To address these issues, we adopt an approach that reduces the overall workload assigned to the LLM and limits its tasks to those that entail semantically informed analysis of the code.
@@
In this paper, we propose the Capture and Resolve Assisted Bounding Strategy (CRABS), a syntactic-semantic hybrid strategy for extracting the information flow graph from a Python notebook. Figure 1 is an overview of the steps employed by CRABS to generate an information flow graph for a given Python notebook. CRABS depends critically on the concept of a notebook’s inter-cell I/O set, defined as the set of all flows of information into or out of individual cells via variables. We first produce lower and upper estimates of the inter-cell I/O set based on the syntactic structure of the notebook, then prompt an LLM to resolve the differences between these two estimates (i.e., ambiguities) one by one. To evaluate this approach, we apply CRABS to 50 representative, highly up-voted Python notebooks from the Meta Kaggle Code dataset (Plotts & Risdal, 2023).
@@ 2 Related Work
LLMs and computational notebooks. Pre-trained language models have been leveraged to generate feedback on answers to open-ended questions about instructional coding or math notebooks (Matelsky et al., 2023); and fine-tuned for code generation (Huang et al., 2022; Yin et al., 2023; Li et al., 2023; Huang et al., 2024). Cao et al. (2024) explore multimodal agents for data science notebook generation.
@@ 3 A Model of Information Flow in Python Notebooks
CRABS views a Python notebook as a sequence of $n$ code cells $C=(c_{1},c_{2},...,c_{n})$ where $c_{j}$ is the $j$th code cell (markdown cells are ignored). CRABS further assumes that cells are executed in order, i.e., from the top to the bottom of the notebook. This assumption limits the possible interpretations of information flows in a notebook; i.e., a value assigned to a variable in cell $c_{j}$ cannot be used in cell $c_{j-1}$ or earlier. We represent the $m$ distinct units of information (i.e., variable values and code references) that flow between cells as $I=\{i_{1},i_{2},...,i_{m}\}$ and define an information flow graph as a set of information flows $F\subseteq\{(c_{s},c_{t},i_{k})\mid 1\leq s<t\leq n\}$, where $i_{k}$ flows from source cell $c_{s}$ to target cell $c_{t}$. The information flow graph in turn implies a cell execution dependency graph, i.e., the set of transitive dependencies $D\subseteq\{(c_{t},c_{s})\mid 1\leq s<t\leq n\}$, where cell $c_{t}$ directly or indirectly depends on cell $c_{s}$. The cell execution dependency graph can be derived from the information flow graph by (1) recognizing a direct dependency between two cells if at least one information flow exists between them in the information flow graph; and (2) computing the transitive closure of this graph to infer indirect dependencies.
@@ 4 Elements of CRABS
Constructing the information flow graph and corresponding cell execution dependency graph for a notebook depends crucially on determining its inter-cell I/O set. The approach taken by CRABS to determine the inter-cell I/O set is based on the observation that some members of the set are certain by shallow syntactic parsing and analysis of the AST, while others are uncertain but may be resolved by taking the semantic meaning of these code cells into consideration. CRABS therefore performs the notebook understanding task in two phases: a syntactic phase and a semantic-aware phase. The syntactic phase employs shallow syntactic parsing to extract an AST based purely on the notebook itself, and analyzes the extracted AST to yield lower and upper estimates of the inter-cell I/O set. The semantic-aware phase resolves the ambiguities between these lower and upper estimates using an LLM.
We represent the inter-cell I/O set as $S\subset\{(c,i_{k},\textit{tag})\mid c\in\{c_{s},c_{t}\},\textit{tag}\in\{\texttt{out},\texttt{in}\}\}$, where the out tag denotes that information $i_{k}$ flows out of source cell $c_{s}$, and the in tag denotes that information $i_{k}$ flows into target cell $c_{t}$. Therefore, it is critical to identify four items for each cell, i.e., data inputs, data outputs, code declarations, and code references. If data $i_{k}$ is defined in cell $c_{s}$ and used in cell $c_{t}$, then $i_{k}$ is a data output of $c_{s}$ and a data input of $c_{t}$. However, when cell $c_{s}$ is examined in isolation, we cannot determine whether there is in fact a subsequent cell $c_{t}$ that uses $i_{k}$. Because we do not provide any information about subsequent cells when asking the LLM to resolve ambiguous information flows associated with a particular cell, we refer to data output candidates rather than data outputs in these prompts. Code references, e.g., functions and classes declared in cells and used in later cells, are treated analogously to data. In the rest of this section, we describe the syntactic phase and semantic-aware phase based on these four items. Relevant examples are attached in the Appendix.
@@ 4.1 Syntactic Phase
The syntactic phase of CRABS performs a limited static analysis of the code in a notebook to yield two estimates of the inter-cell I/O set, representing lower and upper estimates of the actual flows of information into or out of each cell. It produces two estimates because CRABS aims to interpret Python notebooks without accessing the source code for any dependencies, e.g., code defined outside the notebook and accessed via function calls. Deeper or more sophisticated syntactic parsing and analysis of the AST that descends into each function called from a cell likely would achieve tighter lower and upper estimates of the inter-cell I/O set, leaving fewer or even no ambiguities to be resolved by an LLM. However, because we restrict our syntactic analysis to code explicitly included in a notebook’s cells, CRABS does not need to resolve software dependencies. Moreover, even comprehensive static analysis will not yield a unique prediction of the dataflow between cells when conditional control flow structures are employed, e.g., if variables are assigned, used, or updated conditionally (see Appendix A.1.1).
@@
CRABS makes three basic assumptions about the code in Python notebooks relevant to the syntactic phase. First, execution flow follows a strict top-to-bottom order, starting from an initial state with all variables, imports, and other runtime information cleared from memory. Second, all code is reachable. And third, function and class names are not reused as variable names or vice versa (see Appendix A.1.2). In addition, our current implementation has four further limitations. Global variables that are accessed or modified within a function, without being explicitly passed as arguments, will not be recognized on this basis as data inputs or data output candidates by our system (see Appendix A.1.3). Lines of code employing languages other than Python, e.g., shell commands invoked using the ! syntax, or IPython magic commands invoked using the % syntax are ignored (see Appendix A.1.4). A third limitation related to shared data references (aliased variables) is more complex. This issue arises when multiple variables reference the same mutable object in memory, and the object is modified. For example, the first code cell in Figure 5 defines two Pandas DataFrames, train and test. The datasets object is a list that references two mutable objects, train and test. In the second code cell, train is modified in place, resulting in an implicit change in datasets. We define this phenomenon as hidden modifications, wherein a variable name (i.e., datasets) is not explicitly written in a code cell but should still be considered a data output candidate. The current implementation of CRABS detects only those shared references explicitly indicated by assignment (e.g., x = y) or by inclusion in collections (i.e., mutable objects stored in a list, a tuple, or a dictionary). This covers all instances of shared references in our dataset; we will extend CRABS to detect shared references established by other mechanisms in the future. Finally, CRABS produces a single information flow graph per notebook, even when multiple execution paths are possible (e.g., due to conditional statements) given different input data. When different execution paths yield different information flows, the output of CRABS represents the union of those flows.
@@
The lower estimate computed by the syntactic phase represents the set of cell inputs and output candidates considered certain (i.e., unambiguous) given just the code apparent in the cell. This lower estimate is computed using the following rules: (1) Variables explicitly passed to a function are regarded as data inputs only. (2) The iterator of a for loop, a while loop, or a comprehension is regarded as a data input only, unless it is directly reassigned in the loop body. (3) Because CRABS yields information flows for all possible execution paths, it generally treats code in every branch of a conditional statement as contributing to the inputs and the output candidates of the cell. (4) However, when a conditional statement occurs inside a loop, the intersection of all possible execution orders is considered (see Appendix A.1.5 for an illustrative example). (5) The effect of shared references (e.g., aliased variables) is not considered. Figure 6(a) and Figure 6(c) show the information flow graph and the corresponding cell execution dependency graph generated from the lower estimate.
The upper estimate represents the set of cell inputs and output candidates that are either certain or ambiguous, which is a superset of both the ground truth and the lower estimate. This upper estimate is computed using the following rules: (1) Variables explicitly passed to a function are regarded as both data inputs and data output candidates. (2) The iterable of (e.g., the collection traversed by) a for loop, a while loop, or a comprehension is regarded both as a data input and a data output candidate. (3) As with the lower estimate, code on all branches of a conditional statement are generally considered to contribute to the inputs and the output candidates of the cell. (4) But unlike the lower estimate, when a conditional statement occurs inside a loop the union of all possible execution orders is considered. (5) The effect of shared references is taken into account. Figure 6(b) and Figure 6(d) show the information flow graph and the corresponding cell execution dependency graph generated from the upper estimate.
Figure 6: Information flow graphs derived from the (a) lower and (b) upper estimates of the inter-cell I/O set for the notebook in Figure 4; (c) transitive dependencies (solid black) extracted from Figure 6(a) augmented with missing edges (dotted red); and (d) transitive dependencies extracted from Figure 6(b) with correct (black) and incorrect (red) dependencies.
Figures 6(c) and  6(d) demonstrate the subgraph and supergraph relations between the cell execution dependency graphs generated from lower and upper estimates, respectively, and the ground truth. The identified dependencies (black) in Figure 6(c) represent a subgraph of the ground truth, whereas the correct (black) and incorrect (red) edges in Figure 6(d) together represent a supergraph of the ground-truth cell execution dependency graph.